OpenStudy (anonymous):

simplify the radical expression √7 + 3 √7

jimthompson5910 (jim_thompson5910):

hint: think of \(\Large \sqrt{7}\) as x

jimthompson5910 (jim_thompson5910):

if you do that, you have x + 3x

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ \sqrt{7}}}+3{\color{brown}{ \sqrt{7}}}\) as suggested, take common factor

OpenStudy (anonymous):

simplify the radical expression 2√6 + 3√96

OpenStudy (anonymous):

what common factor @jdoe0001

OpenStudy (anonymous):

14sqrt6

OpenStudy (anonymous):

it cant be thats not one of the choices @Shay17

OpenStudy (anonymous):

2sqrt3 + 12sqrt6

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sqrt{96} = \sqrt{16*6}\]

OpenStudy (anonymous):

the choices are 2√7 4√14 4√17 2√14